A car of mass 1250 kg drives around a curve with a speed of 22 m/s.

a. If the radius of the curve is 35 m, what is the centripetal acceleration of the car? (3 points)

b. The maximum friction that the car's tires can apply is 20,000 N. What is the radius of the smallest circle the car can drive in at a speed of 22 m/s? (3 points)


1
Expert's answer
2022-04-27T01:19:05-0400

a) Fc = mv2/r; Fc = 1250 kg • (22 m/c)2/35 m = 785.71

Fc=ma; a = Fc/m = 785.71/1250= 0.629 m/c2

b) Fc = mv2/r; r = mv2/Fc ;

r = (1250 • (22)2)/20000 = 30, 25 m